1. Your hypothesis test for a proportion has a right-tailed critical region. The test statistic is...

1.

Your hypothesis test for a proportion has a right-tailed critical region. The test statistic is zz = 1.84. Find the p-value accurate to 2 decimal places.

p-value =

2.

Your hypothesis test for a mean with n = 23 has a left-tailed critical region. The test statistic is tt = -1.4. Find the p-value accurate to 2 decimal places.

p-value =
Answer #1

solution: right-tailed critical region.

p(z>1.84)=1-p(z<1.84)

=1-0.9671

=0.0329

p value 0.0329

2.

df =n-1 =23-1=22

t =-1.4

The p-value is .087729. using student t table